x over 2 minus 1 equal 8

Answer by richwmiller(17219)

Since you used no parentheses, I will assume that
x over 2 minus 1 equal 8 means
(x/2)-1=8
[It could also mean x/(2-1)=8: x/1=8: x=8]
(x/2)-1=8
add one to both sides
x/2=9
multiply both sides by two
x=18